# Training-free Zero-shot Composed Image Retrieval with Local Concept Reranking

### 14 Dec 2023 v1

@@ 3.2 Local Concept Reranking
Having obtained the positive concepts by parsing modified text in the query, we further determine whether the positive concepts exist for each target image. To detect the local concept existence, we utilize a multimodal model LLaVA [15], which is an instruction-following visual question-answering model, and returns the result by text. However, LLaVA can only detect the existence of local features and return a hard result with the text `Yes' or `No'. This hard result cannot be utilized in the ranking scenario directly. To address this issue, we unlock the last layer of the LLaVA and take the predicted probability of text `Yes' and `No' as the local score, which can be formulated as follows:
$Logit_{YES}^{n},Logit_{NO}^{n}=LLaVA(P_{re},T_{l},I^{n}_{t}),$ (5)
$S_{l}^{n}=softmax(Logit_{YES}^{n},Logit_{NO}^{n}),$ (6)
@@
Given the ranking scores from the first stage of the global retrieval baseline and the second stage of local concept reranking, we can achieve the final ranking score by simply summing the global score and local score. Namely, the final score $S^{n}$ for the $n$-th target image can be shown as follows:
$S^{n}=S_{g}^{n}+S_{l}^{n}.$ (7)
The final ranking is achieved by the $S^{n}$ value from maximum to minimum.
We summarize our proposed method in Algorithm 1.
Algorithm 1 Training-Free Zero-Shot Composed Image Retrieval with Local Concept Reranking.
1:Input: Query of reference image $I_{r}$ and modified text $T_{m}$. Gallery set of target images {$I_{t}\}_{n=1}^{N}$. Hyperparameter rerank top $K$.
@@
6:Local Concept Reranking:
@@
9:     Calculate local score $S_{l}^{n}$ via Eq. (5), Eq. (6).
@@ 4.1 Datasets and Evaluation Metrics
Recall@$K$ Recall@${}_{\text{Subset}}@K$
Training data Method $K=1$ $K=5$ $K=10$ $K=50$ $K=1$ $K=2$ $K=3$
@@
GRB 24.19 52.07 65.48 85.28 60.17 79.13 89.34
GRB + LCR 30.92 56.99 68.58 85.28 66.67 78.68 82.60
Table 1: Quantitative results on CIRR test set. Best scores are highlighted in bold.
We conduct experiments on three publicly available CIR benchmarks, CIRR [18], FashionIQ [29] and CIRCO [1]. All datasets are in the form of triplets, composed of a reference image, modified text, and a target image. FashionIQ, focused on the fashion domain, contains garment images that can be divided into three categories: dress, shirt and toptee. Each image in FashionIQ contains a single subject positioned centrally with a clean background. CIRR is composed of real-life images extracted from NLVR2 [26], which contains rich visual content against various backgrounds. CIRCO, an open-domain dataset, includes 1020 queries divided into 220 for the validation set and 800 for the test set. The dataset utilizes all 120K images of COCO [14] as the gallery set and provides multiple ground truths for each query. Notably, we adopt zero-shot evaluation to test on the test sets of CIRR and CIRCO, and the validation set of FashionIQ following the standard evaluation protocol [25, 1]. The compared conventional methods are fine-tuned on task-specific datasets but our method is training-free. For single ground truth datasets CIRR and FashionIQ, we utilize the top K recall metric for evaluation, and for the multiple ground truth dataset CIRCO, we utilize mean Average Precision (mAP). Specifically, CIRR is evaluated on both the entire gallery and a subset gallery respectively. The subset gallery consists of five highly similar images for each query. FashionIQ evaluation is conducted separately for each category following previous works [25, 1].
@@ 4.2 Implementation Details
We leverage BLIP2 with OPT-2.7b [32] as the default captioner and its image-text matching module with aligned feature space to retrieve target images. For LLM supplying context knowledge, we leverage ChatGPT4-turbo by default to generate pseudo target captions or local discriminative attributes (concepts). For the generated text, we employ a cleaning process to standardize the format and enhance the quality of the text. To achieve scores corresponding to local concepts, we utilize LLaVA-v1.5-13b by default and rerank top 50 target images. All the experiments are conducted on NVIDIA A100 GPUs.
@@
Table 2: Quantitative results on FashionIQ validation set. Best scores are highlighted in bold.
@@
Training-free GRB 23.76 25.31 28.19 29.17
GRB + LCR 25.38 26.93 29.82 30.74
Table 3: Quantitative results on CIRCO test set. Best scores are highlighted in bold.
Figure 3: Qualitative results on CIRR test set. The GRB is the global retrieval baseline, which can merge holistic information from reference image caption and modified text. LCR is local concept reranking, which can extract discriminative local information to rerank based on the results of the global retrieval baseline.
CIRR
Method Captioner LLM R@$1$ R@$5$ R@$10$ R@$50$
GRB BLIP2 GPT4-turbo 24.19 52.07 65.48 85.28
GRB BLIP2 GPT3.5 21.49 45.78 57.11 81.33
GRB instructBLIP GPT4-turbo 22.60 49.88 62.55 84.99
GRB instructBLIP GPT3.5 20.75 45.54 58.10 79.42
Table 4: Ablation studies on CIRR test set. The best scores are highlighted in bold.
CIRR
Method Prompt $P_{l}$ R@$1$ R@$5$ R@$10$
GRB + LCR default 30.92 56.99 68.58
GRB + LCR w/o exemplar 29.54 58.17 69.86
GRB + LCR w/o negative 30.39 58.87 70.27
Table 5: Prompt $P_{l}$ variations with the result on CIRR test set. $P_{l}$ is aimed to parse the modification text and extract local positive concept.
4.3 Comparison with the SOTAs
Competitors.
We compare the proposed TFCIR with several state-of-the-art zero-shot composed image retrieval methods classified into three groups:
  • Training-free methods: We implement three training-free methods for comparison. Text-only-M conducts retrieval directly with the modified text of the triplets. Text-only-C performs retrieval directly with {reference caption} + `change to' + {modified text}, i.e., the concatenation of the modified text and the reference caption of the reference image, which is achieved from the BLIP2 captioner. LLaVA-CIR is achieved by inputting modified text and image, instructing the LLaVA model to output the target caption directly. All of these three methods calculate the cosine similarity distance in the aligned feature space of BLIP2 [13].
4.4 Main Results
CIRR.
The experimental results for the CIRR test set are presented in Table 1. As observed, our baseline method GRB, employing global-level retrieval achieves comparable performance with state-of-the-art ZS-CIR methods SEARLE [1] and Pic2Word [25]. Notably, both of them are trained on text-image pairs, while our method is training-free for saving computational resources. Furthermore, with reranking module LCR incorporating discriminative local concepts, our method gains a significant improvement of 6.72% in Recall@1. Local concepts provide discriminative information, leading to noticeable enhancements. In comparison with three other training-free methods, our method GRB + LCR has an average improvement of 9.01% over Text-only-M, 14.68% over Text-only-C, and 24.49% over LLaVA-CIR respectively. Among these training-free methods, LLaVA-CIR processes detailed information from the reference image, Text-only-C considers the full caption of the reference image, and Text-only-M lacks input from the reference image. It is evident that directly implementing reference image information without careful selection can distract the model and compromise performance. In contrast, the modified text itself provides strong discriminative information and yields competitive results. Our method GRB selectively extracts information from reference image captions by prompt-enhanced language model, resulting in improved performance. Notably, our proposed baseline model can outperform existing training-needed methods [1, 25] by an average improvement of 2.43% over SEARLE and 3.27% over Pic2word. For the evaluation on subset gallery, the discriminative ability is required. Our LCR provides a strong score for the target local concept, increasing the score difference between images, and resulting in a significant improvement of 6.5% over the baseline GRB.
FashionIQ.
We present the results on the FashionIQ validation set in Table 2. We can observe that our GRB baseline alone achieves better performance than the existing state-of-the-art ZS-CIR SEARLE, with an average improvement of 6.81% in Recall@5 and 3.54% in Recall@10. FashionIQ is a fine-grained dataset with similar clean backgrounds and fine-grained variations. We can observe that modified text alone provides strong discriminative information. With the increase of reference image information, the model gets distracted and leads to lower performance from Text-only-C to LLaVA-CIR.
CIRCO.
@@ 4.4 Main Results
As shown in Table 3, we present the performance on the CIRCO test set. As an open domain dataset with multiple ground truths from a much larger gallery set, CIRCO imposes higher requirements of fine-grained discriminative ability. Both of our methods show obvious improvements over existing methods. Specifically, our baseline with global-level retrieval GRB outperforms SEARLE by 12.08% in mAP@5. Additionally, reranking with local concepts LCR further yields over 1% average improvement. The substantial enhancement observed in CIRCO underscores the superiority of addressing reasoning in text over aligning the two modalities.
Visualization Results.
The visualization of the proposed baseline method GRB and reranking method LCR is shown in Figure 3. The baseline method GRB provides coarse direction and the reranking method LCR can identify local discriminative concepts.
@@ 4.5 Ablation Studies
Evaluating Variants of Captioner and LLMs.
In this section, variations of captioner and large language models are identified to provide context knowledge. We implement instructBLIP [11] to generate a descriptive caption by using the prompt `Describe the content of the image in detail'. Additionally, different versions of ChatGPT are deployed to explore the impact of the large language models on the reasoning process. As shown in Table 4, we observe that instructBLIP provides detailed reference information but misleads the model, which is aligned with the observation in Sec. 4.4. Meanwhile, the large language model GPT-4 outperforms GPT3.5, indicative of enhanced reasoning capabilities.
Evaluating Variants of Prompts.
To further evaluate the sensitivity of our model to prompts, we conducted experiments employing diverse prompts, incorporating variations in the number of exemplars and the omission of the negative category. The outcomes of these experiments are summarized in Table 5. Specifically, adhering to the default prompt denoted as $P_{l}$ in Figure 2, we conducted experiments involving the removal of exemplars and the exclusion of the negative category respectively. Under the default prompt $P_{l}$ with manual exemplars and two categories of local concepts, namely positive and negative, the model exhibits superior performance in terms of Recall@1 and a decrease in Recall@5 and Recall@10. In the broader context, our model demonstrates robust performance against prompt variations.
Limitations and Future Work.
From the experimental results, it becomes evident that a judicious selection process is imperative when incorporating information from reference images. In this study, we address this by employing a two-step approach involving initial captioning, followed by the utilization of a Large Language Model (LLM) to generate the pseudo-target caption. The selection of reference information remains an avenue deserving further investigation.
@@ 5 Conclusion
This paper presents an innovative training-free method for zero-shot composed image retrieval, incorporating reranking strategy informed by local concepts. Initially, a global retrieval baseline is established by transforming the text-image composed query into a text-only query. Subsequently, we enhance the retrieval process by extracting local concepts through modified parse techniques and implementing reranking based on the outcomes obtained in the initial stage. The proposed method demonstrates both high efficiency and competitive performance when compared with existing zero-shot composed image retrieval (ZS-CIR) methods.

### 24 Mar 2024 v2

@@ 3.2 Local Concept Re-Ranking
To identify the local concepts that must exist in the target image, we parse the modified text $T_{m}$ by a large language model. We refer to these local concepts that must exist in the target image as positive and those that should not exist as negative. By designing a local concept parsing prompt $P_{l}$ with manually crafted exemplars as shown in Figure 2, we obtain the positive concepts in textual form by removing the negative concepts from the modified text. Notably, it becomes popular to leverage LLMs to generate text descriptions [32, 33, 22] to enhance visual perception. However, current descriptions are typically generated at the class level while our approach generates descriptions at the instance level, specifically tailored to each query. The process can be formulated as follows:
$T_{l}=\text{LLM}(P_{l},T_{m}),$ (4)
where $T_{l}$ denotes the positive local concepts in textual form, providing precise information to enhance discriminative ability.
Having obtained the positive concepts by parsing modified text in the query, we further determine whether the positive concepts exist for each target image candidate. To detect the local concept existence, we utilize a multimodal model LLaVA [17], which is an instruction-following visual question-answering model and returns the result by text. However, LLaVA can only detect the existence of local visual concepts and return a hard result with the text ‘Yes’ or ‘No’, which cannot be utilized in the ranking scenario directly. To address this issue, we unlock the last layer of the LLaVA. Specifically during text generation, LLaVA predicts a score for each token and calculates the next token probability distribution. We focus on the score of tokens ‘Yes’ and ‘No’, and further calculate the visual concept existence probability by exponential normalization, which can be formulated as follows:
$\text{Logit}_{\text{YES}}^{n},\text{Logit}_{\text{NO}}^{n}=\text{LLaVA}(P_{re},T_{l},I^{n}_{t}),$ (5)
$S_{l}^{n}=softmax(\text{Logit}_{\text{YES}}^{n},\text{Logit}_{\text{NO}}^{n}),$ (6)
@@
Given the ranking scores from the first stage of the global retrieval baseline and the second stage of local concept re-ranking, we can achieve the final ranking score by simply summing the global score and local score. Namely, the final score $S^{n}$ for the $n$-th target image can be shown as follows:
$S^{n}=S_{g}^{n}+\gamma\cdot S_{l}^{n}.$ (7)
The final ranking is achieved by the $S^{n}$ value from maximum to minimum. The scale parameter $\gamma$ is a hyperparameter with a default value equal to one. We summarize our proposed method in Algorithm 1.
Algorithm 1 Training-Free Zero-Shot Composed Image Retrieval with Local Concept Re-Ranking.
1:Input: Query of reference image $I_{r}$ and modified text $T_{m}$. Gallery set of target images {$I_{t}\}_{n=1}^{N}$. Hyperparameter rerank top $K$ and scale parameter $\gamma$.
@@
6:Local Concept Re-Ranking:
@@
9:     Calculate local score $S_{l}^{n}$ via Eq. (5) and Eq. (6).
@@ 4.1 Datasets and Evaluation Metrics
We conduct experiments on three publicly available CIR benchmarks, CIRR [20] and CIRCO [1] for scene manipulation, and FashionIQ [31] for fashion attribute manipulation. Additionally, we utilize a general dataset COCO [16] for object composition following [27]. The three CIR benchmarks are in the form of triplets, composed of a reference image, modified text, and one or multiple target images. FashionIQ, focused on the fashion domain, contains garment images that can be divided into three categories: dress, shirt and toptee. CIRR is composed of real-life images extracted from NLVR2 [28], which contains rich visual content against various backgrounds. CIRCO, an open-domain dataset, utilizes all 120K images of COCO [16] as the gallery set and provides multiple ground truths for each query. For object composition following [27], cropped images from COCO validation set (5,000 images) with masked backgrounds are used as queries, with the modified text consisting of a list of object classes. We adopt zero-shot evaluation to test on the test sets of CIRR and CIRCO, and the validation set of FashionIQ following the standard evaluation protocol [27, 1].
Table 1: Quantitative results on CIRR test set. Best scores are in bold.
[…]
Supervision Method $R@1$ $R@5$ $R@10$ $R@50$ $R@1$ $R@2$ $R@3$
@@
GRB 24.19 52.07 64.48 85.28 60.17 79.13 89.34
GRB + LCR 32.82 61.13 71.76 85.28 66.63 78.58 82.68
Figure 3: Qualitative results on CIRR test set. The GRB is the global retrieval baseline, which can merge holistic information from reference image caption and modified text. LCR is local concept re-ranking, which can extract discriminative local information to rerank based on the results of the global retrieval baseline.
@@ 4.2 Implementation Details
We leverage BLIP2 with OPT-2.7b [34] as the captioner and its image-text matching module with aligned feature space to retrieve target images. For LLM supplying context knowledge, we leverage ChatGPT4-turbo to generate pseudo target captions and local discriminative attributes (concepts). For the generated text, we employ a cleaning process to standardize the format and enhance the quality of the text. We set the scale parameter $\gamma$ to 0.1 for the COCO dataset and 1 for all other datasets. To achieve scores corresponding to local concepts, we utilize LLaVA-v1.5-13b and rerank top 50 target images. All the experiments are conducted on NVIDIA A100 GPUs.
4.3 Comparison with the SOTAs
Competitors. We compare the proposed method with several state-of-the-art zero-shot composed image retrieval methods, which are classified into three groups:
  • Training-free methods: We implement three training-free methods for comparison. Text-only-M conducts retrieval directly with the modified text of the triplets. Text-only-C performs retrieval directly with text {reference caption} + ‘change to’ + {modified text}, i.e., the concatenation of the modified text and caption of the reference image, which is achieved from the BLIP2 captioner. LLaVA-CIR is achieved by inputting modified text and image, instructing the LLaVA model to output the target caption directly, with the prompt: ‘Please provide a brief image caption if change to {modified text}. All of these three methods calculate the cosine similarity distance in the aligned feature space of BLIP2 [14].
Table 2: Quantitative results on FashionIQ validation set.
@@ 4.3 Comparison with the SOTAs
GRB + LCR 35.38 55.15 24.84 45.56 33.10 53.60 31.11 51.44
4.4 Main Results
CIRR. The experimental results for the CIRR test set are presented in Table 1 and visualization results are shown in Figure 3. As observed, our global retrieval baseline (GRB) achieves comparable performance with state-of-the-art ZS-CIR methods SEARLE [1] and Pic2Word [27]. Notably, both of them are trained on text-image pairs, while our method is training-free for saving computational resources. Furthermore, with local concepts re-ranking (LCR) incorporating discriminative information, our method gains a significant improvement of 8.63% in Recall@1. In comparison with three other training-free methods, our method GRB + LCR has an average improvement of 11.32% over Text-only-M, 16.99% over Text-only-C, and 24.80% over LLaVA-CIR respectively. To process modification reasoning, LLaVA-CIR leverages a vision-language model and Text-only-C relies on simple combination. However, it is evident that existing text-image alignment space (BLIP2) cannot handle instructional reasoning effectively, and relying on vision-language model (LLaVA) only leads to suboptimal results. Our GRB addresses modification reasoning by transforming the instructions into descriptive pseudo-target captions using a prompt-enhanced pure language model, resulting in improved performance. Notably, our proposed baseline model can outperform existing training-needed methods [1, 27] by an average improvement of 4.74% over SEARLE and 5.57% over Pic2word.
FashionIQ. We present the results on the FashionIQ validation set in Table 2. We can observe that our GRB baseline alone achieves better performance than the existing state-of-the-art ZS-CIR SEARLE, with an average improvement of 6.81% in Recall@5 and 3.54% in Recall@10. FashionIQ is a fine-grained dataset with similar clean backgrounds and fine-grained variations. We can observe that modified text alone provides strong discriminative information. With the increase of reference image information, the model gets distracted and leads to lower performance from Text-only-C to LLaVA-CIR.
Table 3: Quantitative results on CIRCO test set. Best scores are in bold.
@@ 4.4 Main Results
Training-free Text-only-M 4.40 4.81 5.59 5.99
Text-only-C 15.28 16.45 18.48 19.39
LLaVA-CIR 12.13 12.88 14.45 15.2
GRB 23.76 25.31 28.19 29.17
GRB + LCR 26.52 28.25 31.23 31.99
@@
CIRCO. As shown in Table 3, we present the performance on the CIRCO test set. As an open domain dataset with multiple ground truths from a much larger gallery set, CIRCO imposes higher requirements of fine-grained discriminative ability. Both of our methods show obvious improvements over existing methods. Specifically, our baseline with global-level retrieval GRB outperforms SEARLE by 12.08% in mAP@5. Additionally, re-ranking with local concepts LCR further yields 2.89% average improvement. The substantial enhancement observed in CIRCO underscores the superiority of addressing reasoning in text over aligning the two modalities.
COCO. Table 4 and Figure 4 present the quantitative and qualitative performance on COCO [16] for object composition following [27], respectively. Note that the modified text is simplified as the name of classes, thus our GRB is
Table 4: Quantitative result on COCO.
Supervision Method R@1 R@5 R@10
Text-image pairs Pic2word 11.5 24.8 33.4
Training-free GRB 15.6 30.7 38.8
GRB+LCR 29.2 55.6 69.8
equivalent to Text-only-C in this scenario. With LCR, our method achieves a significant 17.7% improvement in R@1 over the existing state-of-the-art method Pic2Word.
Figure 4: Qualitative results on COCO for object composition.
@@ 4.5 Ablation Studies
Evaluating Variants of Captioner and LLMs. In this section, as shown in Table 5, we evaluate various captioners and large language models for the global retrieval baseline. We implement instructBLIP [13] to generate a detailed caption by employing the prompt ‘Describe the content of the image in detail’, contrasting with BLIP2, which provides a concise caption. Furthermore, different versions of ChatGPT are deployed to explore the impact of large language models on the reasoning process. We observe that instructBLIP provides detailed reference information but can mislead the model. Meanwhile, the large language model GPT4 outperforms GPT3.5, indicating enhanced reasoning capabilities.
Evaluating Variants of Baselines. Figure 4(a) compares LCR with additional baselines including training-free methods Image-only, Text-only-M, and a triplet-supervised method CLIP4Cir [2]. Our LCR shows consistent and significant improvement across both training-free and training-based baselines.
Evaluating Variants of Re-rank Top $K$. Figure 4(b) compares LCR with various top $K$. We observed that the improvement is significant initially and stabilizes with larger $K$, indicating that a small effort (re-rank top 5) in the re-ranking process can yield noticeable improvements (7.2% in R@1).
Evaluating Variants of Prompts $P_{l}$. Figure 4(c) evaluates the sensitivity of our model to prompts $P_{l}$. Similar to the default prompt denoted as $P_{l}$ in Figure 2, Prompt1 is without examplars and instructs the model to return positive and negative concepts by noun phrases. Prompt2 is with additional exmplars and returns the output also by noun phrase. Prompt 3, our default prompt, is with additional examplars and instructs the model to return sentences. We observe that both detailed information and additional examplars can bring improvement.
(a) Baseline Variations
(b) Re-rank Top $K$ Variations
(c) Prompt Variations
Figure 5: Ablation study on CIRR test set.
Table 5: Ablation study on CIRR and CIRCO test set.
CIRR Recall CIRCO mAP@$k$
Captioner LLM R@$1$ R@$5$ R@$10$ R@$50$ $k=5$ $k=10$ $k=25$ $k=50$
InstructBLIP GPT3.5 20.75 45.54 58.10 79.42 11.52 12.36 14.19 14.92
InstructBLIP GPT4 22.60 49.88 62.55 84.99 16.46 17.57 19.68 20.56
BLIP2 GPT3.5 21.49 45.78 57.11 81.33 21.32 22.81 25.25 26.27
BLIP2 GPT4 24.19 52.07 64.48 85.28 23.76 25.31 28.19 29.17
Limitations and Future Work. From the experimental results, it becomes evident that a careful construction of query information is imperative, especially to determine the requirement that must be satisfied. However, the requirement supplied by the reference image is ambiguous. In this study, we address this by re-ranking on the local concepts that must be satisfied. The conditional selection of reference image information remains an avenue deserving further investigation.
@@ 5 Conclusion
This paper presents an innovative training-free method for zero-shot composed image retrieval, incorporating re-ranking strategy informed by discriminative local concepts. Global retrieval baseline converts the instruction-image query into a descriptive caption to leverage existing visual-textual alignment space. The second stage, local concept re-ranking, solves the problem of ambiguous requirements from the reference image by re-ranking on precise fine-grained concepts. The proposed method demonstrates comparable performances to that of existing training based zero-shot composed image retrieval methods and outperforms significantly all other training-free methods.